Suppose 40 percent of the customers at Pizza Palooza order a square pizza, 80 percent order a soft drink, and 32 percent order both a square pizza and a soft drink.

Is ordering a soft drink independent of ordering a square pizza?
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

2 events are independent if P(A and B) = P(A)*P(B)

P(A and B) in our case is .32

P(A) = .40

P(B) = .80

P(A)*P(B) = .40*.80 = .32

Now, P(A and B) = P(A)*P(B): so A and B are independent.

Ordering a soft drink IS independent of ordering a square pizza.